OpenStudy (anonymous):

help please...logx-2^125=3

OpenStudy (anonymous):



OpenStudy (anonymous):

I need to solve for x

OpenStudy (anonymous):

\[\log_{x-2}(125)=3\]?

OpenStudy (anonymous):

this means \[(x-2)^3=125\]

OpenStudy (anonymous):

which tells you \(x-2=5\)

OpenStudy (anonymous):

yes I went up to there...so then I subtracted 2 from 5

OpenStudy (anonymous):

wait nvrm what i said there